OpenStudy (anonymous):

Raya settled on a price of $5,225 for a subcompact car. The dealer had to add 6% sales tax. Raya had no car trade in, but she wanted to make a 25% down payment. Raya then decided to repay the remainder of the car off in 36 monthly payments. Her monthly payments, which included the finance charge, were $127.50. What was the final cost of the new car?

OpenStudy (mayankdevnani):

price of car times 6% tax: 5225 x .06 = 313.50 tax price plus tax = 5538.50 down payment of 25%: 5538.50 x .25 = 1384.63 down remainder of balance is 4153.87/ 36 months to pay = 115.39 as payment with no finance charge payments are 127.50 - 115.39 = 12.11 in finance charges monthly 12.11 x 36 = 435.96 total for car and tax was 5538.50 + 435.96 in finance charges = $5974.46 total paid for car http://answers.yahoo.com/question/index?qid=20120215114037AAswQJe

OpenStudy (anonymous):

thank you so much